# Early-Cycle Internal Impedance Enables ML-Based Battery Cycle Life Predictions Across Manufacturers

### v1

@@
Early-Cycle Internal Impedance Enables ML-Based Battery Cycle Life Predictions Across Manufacturers
@@
Predicting the end-of-life (EOL) of lithium-ion batteries across different manufacturers presents significant challenges due to variations in electrode materials, manufacturing processes, cell formats, and a lack of generally available data. Methods that construct features solely on voltage-capacity profile data typically fail to generalize across cell chemistries. This study introduces a methodology that combines traditional voltage-capacity features with Direct Current Internal Resistance (DCIR) measurements, enabling more accurate and generalizable EOL predictions. The use of early-cycle DCIR data captures critical degradation mechanisms related to internal resistance growth, enhancing model robustness. Models are shown to successfully predict the number of cycles to EOL for unseen manufacturers of varied electrode composition with a mean absolute error (MAE) of 150 cycles. This cross-manufacturer generalizability reduces the need for extensive new data collection and retraining, enabling manufacturers to optimize new battery designs using existing datasets. Additionally, a novel DCIR-compatible dataset is released as part of ongoing efforts to enrich the growing ecosystem of cycling data and accelerate battery materials development.
@@ 1 Introduction
Lithium-ion batteries are indispensable across a wide range of applications due to their high energy densities, decreasing costs, and extended lifetimes. However, predicting the end of life (EOL) of these batteries, which is critical to ensuring their reliability and safety, remains a significant challenge due to the complex and nonlinear degradation mechanisms involved. Early and accurate prediction of battery EOL is essential to enable accelerated testing and validation of new battery formulations, allowing manufacturers to optimize chemistries and processes more efficiently.
@@
Severson et al. (2019) made significant contributions by utilizing early-cycle features from standard discharge capacity curves to predict battery life, introducing an approach that provided high accuracy for end-of-life (EOL) predictions. Since then, a wide array of studies has focused on improving EOL predictions (Elmahallawy et al., 2022; Ling, 2022; Paulson et al., 2022; Saxena et al., 2022; Ma et al., 2020; Fei et al., 2022; Geslin et al., 2023; Saxena et al., 2021; Chen et al., 2022; Li et al., 2019; Rangel-Martinez et al., 2021; Saxena et al., 2015; Zhang et al., 2018; Hosen et al., 2021; Attia et al., 2021; Ng et al., 2020). However, many of these works have relied on capacity-based features and models that perform well only within the specific datasets on which they were trained. When applied to different chemistries or operating conditions, their predictive power diminishes significantly (Severson et al., 2019; Sulzer et al., 2021; Finegan et al., 2021; Attia et al., 2020; Paulson et al., 2024; Finegan and Cooper, 2019; Walker et al., 2019; Finegan et al., 2019; Wang et al., 2024). (Paulson et al., 2022) created a machine learning model that generalized across several cathode chemistries, utilizing a dataset drawn from a substantial internal archive of batteries produced in-house. Recently, (Rahmanian et al., 2024) proposed an attention-based recurrent neural network capable of handling diverse chemistries and cycling protocols. While their model demonstrates adaptability, it still requires fine-tuning for new chemistries and is reliant on large proprietary datasets, limiting its practical application in environments with limited data availability. Therefore, predictive models for battery cycle life that generalize across manufacturers without relying on data-hungry models are still lacking. Battery performance is influenced not only by electrode chemistry but also by a variety of factors related to cell design and manufacturing processes (Baumhöfer et al., 2014). Even cells with nominally similar chemistries can exhibit vastly different behaviors due to differences in manufacturing parameters, including electrode coating thickness, particle size distribution, slurry preparation, and precision in assembly. Such variations can affect the electrochemical performance, aging mechanisms, and overall reliability of a cell, making cross-manufacturer generalization a significantly more complex challenge than cross-chemistry generalization alone. These subtle changes profoundly impact degradation pathways, leading to different rates and patterns of failure. For instance, as a battery cell cycles, degradation processes such as the formation and growth of the solid electrolyte interphase (SEI) on the anode surface, lithium plating, and the accumulation of inactive lithium deposits lead to an increase in interfacial resistance. Concurrently, mechanical degradation of electrode materials, loss of active material, and electrolyte decomposition hinder ion and electron mobility, resulting in an overall increase in internal impedance (Diao et al., 2022; Fermín-Cueto et al., 2020).
@@
To address this challenge, the present work uses Direct Current Internal Resistance (DCIR) measurements from early cycles as a key feature in predictive modeling, capturing underlying degradation processes to enable accurate EOL predictions across cell manufacturers and operating conditions. Using the ensemble of DCIR and traditional capacity-based features from early cycles, models are developed that maintain high predictive accuracy even with simple linear regressors, achieving intrinsic generalizability without the need for complex algorithms or extensive retraining. Incorporating DCIR into routine cycling protocols offers a straightforward and cost-effective way to enhance model robustness and generalizability across varying manufacturers.
@@ 2 Cycling Protocol and Dataset Generation
Figure 1: Distinct cycling behavior across battery manufacturers complicates cross-manufacturer prediction when relying on capacity-based features alone. (a) Discharge capacity as a function of cycle number for batteries from three manufacturers (LG MH1, Samsung 50E, and BAK CG). Clear differences in degradation rates and capacity loss profiles can be observed across manufacturers. (b-d) Differential capacity (dQ/dV) curves at cycle 100 for (b) BAK CG, (c) Samsung 50E, and (d) LG MH1 cells, highlighting the distinct electrochemical signatures for each manufacturer. (e) t-SNE plot of features derived from cycling data, showing clustering of cells by manufacturer, further illustrating the challenge of cross-manufacturer generalization in battery life prediction.
To address the scarcity of standardized cross-chemistry datasets for battery end-of-life (EOL) prediction, a comprehensive benchmark dataset acquired from 57 commercial lithium-ion cells from 3 different manufacturers with different cathode-anode combinations is introduced (Fig.1). The cells were cycled with a diagnostic protocol that incorporated periodic DCIR measurements to track the internal resistance of the battery cells at incremental states-of-charge (SOC) to provide a detailed view of cell degradation and performance, a method widely utilized by OEMs and cell makers.
Battery cycling data were recorded at intervals ranging from a few seconds to a few minutes, depending on the cycling current, generating time-series data for voltage, current, and time. From these fundamental quantities, additional metrics such as capacity and energy were computed. A key metric, the discharge capacity, was tracked across cycles to monitor battery health, with the remaining percentage of initial discharge capacity serving as the state-of-health (SOH).
The discharge capacity versus cycle number for all cells is illustrated in Fig. 1a, where cells are colored by manufacturer. While most cells exhibit linear decay in discharge capacity, some demonstrate non-linear profiles, emphasizing the need for models that can accurately capture these variations. The focus of this work is on predicting the cycle number at which a fixed SOH (e.g., 85%) is reached, recognizing that different applications may require different SOH thresholds for a cell to be considered at its ”end of life”. With this in mind, it should be noted that while ”EOL” may not be the most appropriate terminology for this predictive task, it is used throughout the manuscript for consistency with other works in the field.
@@ 2.1 Battery cell manufacturers and chemistries
The dataset presented in this work is comprised of 57 Li-ion battery cells from 3 different manufacturers, each with a distinct cathode-anode combination. Table 1 describes the difference between each cell type. The cathode in all cells are considered high-Ni; the Samsung cells contain a Nickel-Cobalt-Aluminum (NCA) cathode (Popp et al., 2020; Schmitt et al., 2023) while the BAK and LG cells contain a Nickel-Manganese-Cobalt (NMC) cathode. The anode in all cells is graphite-based; the Samsung and BAK cells contain a Graphite-Silicon (Gr-Si) blend anode, while the LG cells contain a pure Graphite (Gr) anode.
@@
These cells differ in their electrode compositions and form factors, which directly influence their degradation behaviors. Fig.1a shows that the discharge capacity versus cycle number curves have varying characteristics depending on the manufacturer and cycling conditions; linear, sub-linear, and super-linear decay profiles are observed (Attia et al., 2022). Fig.1b-d demonstrates the differences in electrochemistry between the three cell types considered through the evolution of the differential capacity profiles over 50 cycles (data collected on a high-resolution NOVONIX UHPC system). To visualize the differences in aging characteristics between each cell type, the differential capacities are used in a t-distributed stochastic neighbor embedding, shown in Fig.1e.
By incorporating this diversity, the presented dataset is designed to validate the generalizability of machine learning models across varying commercial cell chemistries. This approach ensures that findings presented in this work are not limited to a single cell type but are applicable across a broad spectrum of commercially relevant lithium-ion cells.
@@ 2.2 Operating conditions
Table 2 summarizes the various cycling conditions comprising this dataset. The cells were tested across three voltage ranges and three temperatures to capture a broad spectrum of degradation behaviors. Notably, only the Sam50E and BakCG cells were cycled within a 3.0 V - 4.2 V range to avoid Li-Si alloying reactions, which are specific to Gr-Si anodes. Cycling at 60°C was conducted exclusively with the 2.5 V - 4.2 V range because at high-temperatures degradation is expected to be dominated by time-dependent reactions rather than cycle-dependant reactions.
@@
T V 2.5 - 4.06 2.5 - 4.2 3.0 - 4.2
25 Sam50E, BakCG, LGMH1 Sam50E, BakCG, LGMH1 Sam50E, BakCG
40 Sam50E, BakCG, LGMH1 Sam50E, BakCG, LGMH1 Sam50E, BakCG
60 - Sam50E, BakCG, LGMH1 -
Table 2: Battery cell cycling conditions matrix, where T refers to the temperature (in Celsius) and V refers to the volatge range between which the cells were cycled. Three voltage ranges and three temperatures were employed to cover various degradation modes across three different cell manufacturers.
@@ 40 Sam50E, BakCG, LGMH1 Sam50E, BakCG, LGMH1 Sam50E, BakCG
This protocol efficiently captures dynamic performance, internal resistance, and overall battery health under realistic operating conditions, providing a rich dataset for developing predictive models that can generalize across various chemistries and conditions.
@@ 3 Machine Learning Approach
Figure 2: Schematic representation of feature extraction for battery life prediction. Starting with raw discharge capacity curves, features such as $\Delta Q(V)$, linear capacity fade, and DCIR-based $\Delta Q(V)$ features are extracted for state-of-health prediction.
All features were extracted from time-series data of the first 102 cycles (Fig.2). Emphasis was given to features derived from electrochemically meaningful data fields such as internal resistance, calculated from Direct Current Internal Resistance (DCIR) pulses:
@@
    $\Delta Q(V)$ Features — Following the work of (Severson et al., 2019), univariate time-series features were extracted from the difference between discharge capacity vs. potential curves, $Q(V)$, for two specific cycles (in this case, cycles 101 and 1, corresponding to the slower C/10 rate step prior to the DCIR checkup cycle). The extracted features include skewness ($\gamma_{1}$), kurtosis ($\gamma_{2}$), variance ($\sigma^{2}$), mean ($\mu$), minimum ($\min$), and maximum ($\max$) of the function:
@@
    where $Q_{n}(V)$ represents the $Q(V)$ curve at cycle $n$. These statistical features capture the distributional and shape changes in $Q(V)$ curves, which provide insights into the battery’s state of health over time. Fig. 2 depicts sample $Q_{1}(V)$ and $Q_{101}(V)$ curves, with the inset showing the corresponding $\Delta Q_{101-1}(V)$.
@@
    $R_{\text{DCIR}}$ Features — Each DCIR measurement involves applying a current pulse at a specific state of charge (SOC) during a discharge step. The response to a current pulse is a voltage drop that allows for the calculation of an effective internal resistance: $R=V/I$. To build features for machine learning models, resistance values were extracted for each of the 10 current pulses during cycles 2 and 102 (Fig. 2).
@@
    $\Delta Q_{\text{DCIR}}(V)$ Features — Recognizing that each current pulse and its associated voltage drop response from DCIR measurements can be viewed as forming a partial $Q(V)$ curve, the methodology from (Severson et al., 2019) is extended to calculate the same set of univariate time-series statistical transformations (mean, minimum, maximum, variance, skewness, and kurtosis) on these partial curves across the 10 SOC levels, between cycles 2 and 102. This process generated 60 additional features, which is referred to as $\Delta Q^{(k)}_{\text{DCIR}}(V)$ features, where $k$ represents the pulse number. This approach enables us to capture the nuances in the capacity changes at various SOC levels, providing a more detailed understanding of the degradation process. For instance, the kurtosis of the changes in these partial capacity curves ($\gamma_{2}(\Delta Q^{(k)}_{\text{DCIR},102-2}(V))$) can reveal sharpness or outlier behavior in capacity changes, which might correlate with specific degradation mechanisms.
@@
    Linear Fit Features — To capture the overall trend in discharge capacity over time, a linear fit is applied to the discharge capacity vs. cycle number curve between two early cycles (e.g., cycles 30 and 99). The slope ($p_{1}$) and intercept ($p_{2}$) of this linear fit were used as features. These features provide insights into the rate of capacity fade and the initial state of the battery, which are crucial for predicting future performance. Although more complex fits were explored, they did not significantly improve model performance and were therefore not included in the final feature set.
@@
These features were then combined and used to train various machine learning models. For regression-based modeling, the best performance was achieved using a simple linear elastic net model. While more complex models such as Random Forest and Multilayer Perceptron were tested (as detailed in the Appendix), Elastic Net was ultimately selected due to its regularization properties, which help mitigate overfitting—especially when using a relatively small dataset or when the features exhibit multicollinearity. This is particularly important given the nature of our feature sets, which include correlated variables from both traditional capacity-based features and those derived from DCIR measurements. Additionally, as a preprocessing step, Sequential Feature Selection (SFS) was applied for high dimensional feature sets (e.g., the DCIR features) to refine the input feature space for subsequent Principal Component Analysis (PCA). This method demonstrates the robustness of the selected features which achieve high prediction accuracy without the need for complex or resource-intensive models. Although the models trained in this work show strong predictive accuracy, it is noted that a larger and more diverse dataset is essential to further validate these findings and ensure broader applicability. While cross-validation provides a robust estimate of model performance, the addition of an independent holdout dataset would offer a more definitive evaluation of the model’s capacity to generalize across different chemistries, manufacturers, and operating conditions.

### v2

# Early-Cycle Current Pulses Enable ML-Based Battery Cycle Life Predictions Across Manufacturers

@@
Early-Cycle Current Pulses Enable ML-Based Battery Cycle Life Predictions Across Manufacturers
@@
Predicting the end-of-life (EOL) of lithium-ion batteries across different manufacturers and operating conditions presents significant challenges due to variations in electrode materials, electrolyte formulations, manufacturing processes, cell formats, and a lack of publicly available data. Methods that construct features solely on voltage-capacity profile data typically fail to generalize across cell chemistries. This study introduces a novel methodology to develop features using Direct Current Internal Resistance (DCIR) measurements from current pulses that span the available state-of-charge range within the first 100 cycles of cell life, enabling more accurate and generalizable EOL predictions. The use of early-cycle DCIR data captures critical degradation mechanisms and enhances model robustness. Models are shown to successfully predict the number of cycles to EOL, ranging from 700 to 2400 cycles, for unseen cell types containing varied electrode compositions with mean absolute errors (MAE) of 105, 160, and 230 cycles for EOL conditions of 90, 85, and 80 %, respectively. This generalizability between manufacturers reduces the need for extensive new data collection and retraining, enabling manufacturers to optimize new battery designs using existing datasets and short-term testing. Additionally, a novel DCIR-compatible dataset is released as part of ongoing efforts to enrich the growing ecosystem of cycling data and accelerate battery materials development.
@@ 1 Introduction
Lithium-ion batteries are indispensable energy storage devices across a wide range of applications due to their high energy densities, decreasing costs, and extended lifetimes. However, predicting the end of life (EOL) of these batteries based on capacity loss, which is critical to ensuring their reliability and safety, remains a significant challenge due to the complex and nonlinear degradation mechanisms involved. Early and accurate prediction of battery EOL is essential to enable accelerated testing and validation of new battery formulations, allowing manufacturers to optimize chemistries and processes more efficiently.
@@
Severson et al. (2019) demonstrated that by utilizing early-cycle features from constant-current voltage versus discharge capacity curves, battery life could be predicted with high accuracy within a single cell chemistry. Since then, a wide array of studies have focused on improving EOL predictions (Elmahallawy et al., 2022; Ling, 2022; Paulson et al., 2022; Saxena et al., 2022; Ma et al., 2020; Fei et al., 2022; Geslin et al., 2023; Saxena et al., 2021; Chen et al., 2022; Li et al., 2019; Rangel-Martinez et al., 2021; Saxena et al., 2015; Zhang et al., 2018; Hosen et al., 2021; Attia et al., 2021; Ng et al., 2020). However, many of these works have relied on features derived from constant-current voltage versus capacity curves, leading to models that perform well only within the specific datasets on which they were trained. When applied to different chemistries or operating conditions, their predictive power diminishes significantly (Severson et al., 2019; Sulzer et al., 2021; Finegan et al., 2021; Attia et al., 2020; Paulson et al., 2024; Finegan and Cooper, 2019; Walker et al., 2019; Finegan et al., 2019; Wang et al., 2024). Paulson et al. (2022) created a machine learning model that generalized across several cathode chemistries, utilizing a dataset drawn from a substantial internal archive of batteries produced in-house. Recently, Rahmanian et al. (2024) proposed an attention-based recurrent neural network capable of handling diverse chemistries and cycling protocols. While their model demonstrates adaptability, it requires fine-tuning for new chemistries and is reliant on large proprietary datasets. This reliance could potentially limit its practical application in environments with restricted data availability and application to new datasets. For these reasons, there remains a need for battery cycle life prediction models that generalize across manufacturers and operating conditions without relying on data-hungry models. To add further complexity to the problem, battery performance is influenced not only by electrode chemistry and electrolyte formulation but also by a variety of factors related to cell design and manufacturing processes (Baumhöfer et al., 2014). Even cells with nominally similar chemistries can exhibit vastly different behaviors due to differences in manufacturing, including electrode coating thickness, particle size distribution, slurry preparation, and precision in assembly. Such variations can affect the electrochemical performance, aging mechanisms, and overall reliability of a cell, making cross-manufacturer generalization a significantly more complex challenge than cross-chemistry generalization alone. These subtle changes profoundly impact degradation pathways, leading to different rates and patterns of failure. For instance, as a battery cell cycles, degradation processes such as the formation and growth of the solid electrolyte interphase (SEI) on the anode surface, lithium plating, and the accumulation of inactive lithium deposits lead to an increase in interfacial resistance. Concurrently, mechanical degradation of electrode materials, loss of active material, and electrolyte consumption hinder ion and electron mobility, resulting in an overall increase in internal impedance (Diao et al., 2022; Fermín-Cueto et al., 2020).
@@
To address this challenge, the present work employs current pulses across the state-of-charge range, known as Direct Current Internal Resistance (DCIR) measurements, from early cycles to derive key features for predictive modeling. This approach effectively captures underlying degradation processes to enable accurate EOL predictions across cell manufacturers and operating conditions. Using the ensemble of DCIR and traditional voltage-capacity curve features from early cycles, models are developed that maintain high predictive accuracy even with simple linear regressors, achieving intrinsic generalizability without the need for complex algorithms or extensive retraining. Feature selection demonstrates the importance of current pulses in early life to obtain generalizable predictions. Incorporating DCIR measurements into routine cycling protocols offers a straightforward and cost-effective way to enhance model robustness and generalizability across varying manufacturers.
@@ 2 Cycling Protocol and Dataset Generation
Figure 1: Distinct cycling behavior across battery manufacturers complicates cross-manufacturer prediction. (a) Discharge capacity as a function of cycle number for battery cells from three manufacturers (LG MH1, Samsung 50E, and BAK CG) under varied operating conditions. Clear differences in degradation rates and capacity loss profiles can be observed across manufacturers. (b-d) Differential capacity (dQ/dV) curves of the first 50 cycles for (b) BAK CG, (c) Samsung 50E, and (d) LG MH1 cells, highlighting the distinct electrochemical signatures of each cell type. (e) t-SNE plot of features derived from cycling data, showing clustering of cells by manufacturer, illustrating the challenge of model generalization across cell types in battery lifetime predictions.
To address the scarcity of standardized datasets across multiple cell chemistries, a comprehensive benchmark dataset acquired from 57 commercial lithium-ion cells from 3 different manufacturers, each with a different cathode-anode combination, is introduced (Fig. 1). The cells were cycled with a diagnostic protocol that incorporated periodic direct-current internal resistance (DCIR) measurements to track the internal resistance of the battery cells in incremental states-of-charge (SOC) to provide a detailed view of cell degradation and performance, a method widely utilized by OEMs and cell makers.
Battery cycling data were recorded at intervals ranging from a few seconds to a few minutes, depending on the cycling current, generating time-series data for voltage, current, and time. From these fundamental quantities, additional metrics such as capacity and energy were computed. A key metric, the discharge capacity, was tracked across cycles to monitor battery health, where the remaining percentage of initial discharge capacity is referred to as the battery cell’s state-of-health (SOH).
The discharge capacity versus cycle number for all cells is illustrated in Fig. 1a, where cells are colored by manufacturer. While most cells exhibit linear decay in discharge capacity, some demonstrate non-linear profiles, emphasizing the need for models that can accurately capture these variations. The focus of this work is on predicting the cycle number at which a fixed SOH (e.g., 85%) is reached, recognizing that different applications may require different SOH thresholds for a cell to be considered at its ”end of life”. With this in mind, it is noted that while ”EOL” may not be the most appropriate terminology for this predictive task, it is used throughout the manuscript for consistency with other works in the field. In this regard, ”EOL” refers to a target SOH.
@@ 2.1 Battery cell manufacturers and chemistries
The dataset presented in this work is comprised of 57 Li-ion battery cells from 3 different manufacturers, each with a distinct cathode-anode combination. Table 1 describes the difference between each cell type. The cathode in all cells are considered high-Ni; the Samsung INR21700-50E (Sam-50E) cells contain a Nickel-Cobalt-Aluminum (NCA) cathode (Popp et al., 2020; Schmitt et al., 2023) while the BAK N21700CG (BAK-CG) and LG INR18650-MH1 (LG-MH1) cells contain a Nickel-Manganese-Cobalt (NMC) cathode. The anode in all cells is graphite-based; the Sam-50E and BAK-CG cells contain a Graphite-Silicon (Gr-Si) blend anode, while the LG-MH1 cells contain Graphite-only (Gr) anode.
@@
These cells differ in their electrode compositions and form factors, which directly influence their degradation behaviors. Fig. 1a shows that the discharge capacity versus cycle number curves have varying characteristics depending on the manufacturer and cycling conditions; linear, sub-linear, and super-linear decay profiles are observed (Attia et al., 2022). Fig. 1b-d demonstrates the differences in electrochemistry between the three cell types, shown by the evolution of the differential capacity profiles over 50 cycles (data collected on a high-resolution NOVONIX UHPC system at 40 °C and a rate of C/10). To visualize the differences in aging characteristics between each cell type, the differential capacities are used for a t-distributed stochastic neighbor embedding (t-SNE), shown in Fig. 1e.
By incorporating this diversity, the presented dataset is designed to validate the generalizability of machine learning models across commercial cell chemistries. This approach ensures that the findings presented in this work are not limited to a single cell type but are applicable across a broad spectrum of commercially relevant lithium ion cells.
@@ 2.2 Operating conditions
Table 2 summarizes the various cycling conditions that comprise this dataset. Cells were tested within three voltage ranges and at three temperatures to capture a broad spectrum of degradation behaviors. The 3.3 - 4.2 V range was chosen to avoid accessing capacity from the Si component of the Gr-Si negative electrodes, which occurs at lower voltages where Li alloys with Si. For this reason, the Sam-50E and BAK-CG cells were cycled within a 3.3 V - 4.2 V range but the LG-MH1 cells were not because they do not contain any Si. Cycling at 60 °C was conducted exclusively within the 2.5 V - 4.2 V range.
@@
[…]
2.5–4.06 2.5–4.2 3.3–4.2
23 Sam-50E, BAK-CG, LG-MH1 Sam-50E, BAK-CG, LG-MH1 Sam-50E, BAK-CG
40 Sam-50E, BAK-CG, LG-MH1 Sam-50E, BAK-CG, LG-MH1 Sam-50E, BAK-CG
60 Sam-50E, LG-MH1
Table 2: Cell testing conditions used in this study, grouped by temperature and voltage range.
@@ 60 Sam-50E, LG-MH1
This protocol efficiently captures dynamic performance and overall battery health under realistic operating conditions, providing a rich dataset for developing predictive models that can generalize across various chemistries and conditions.
@@ 3 Machine Learning Approach
Figure 2: Schematic representation of feature extraction for battery cycle life prediction. Starting with 102 cycles of measured time-series data for different cell types and operating conditions, features are computed from differences in voltage-capacity profiles using both the full voltage range and current pulses, and effective resistances (DCIR). The feature set dimensions are reduced using sequential feature selection (SFS) and principle component analysis (PCA). Models are trained to predict the number of cycles to a target end-of-life EOL condition.
3.1 Feature set
Comprehensive feature sets were extracted from time-series data of the first 102 cycles (Fig. 2). The feature sets were down-selected before assessing model performance, and section 4.2 presents associated implications, particularly how selected features preferentially correspond to electrochemically meaningful quantities. The range of features considered in this work are:
@@ 3.1 Feature set
    $\Delta Q(V)$ Features — Following the work of Severson et al. (2019), univariate time-series features were extracted from the difference between discharge capacity vs. potential curves, $Q(V)$, for two specific cycles (in this case, cycles 101 and 1, corresponding to the slower C/10 rate step prior to the DCIR checkup cycle). The extracted features include skewness ($\gamma_{1}$), kurtosis ($\gamma_{2}$), variance ($\sigma^{2}$), mean ($\mu$), minimum ($\min$), and maximum ($\max$) of the function:
@@
    where $Q_{n}(V)$ represents the $Q(V)$ curve at cycle $n$. These statistical features capture the distributional and shape changes in $Q(V)$ curves. Fig. 2 depicts sample $Q_{1}(V)$ and $Q_{101}(V)$ curves, with the inset showing the corresponding $\Delta Q_{101-1}(V)$.
@@
    $R_{\text{DCIR}}$ Features — Each DCIR measurement involves applying a current pulse at a specific state of charge (SOC) during a discharge step. The response to a current pulse is a voltage drop that allows calculating an effective internal resistance: $R=\Delta V/I$. Included in the feature sets were resistance values extracted for each of the 10 current pulses during cycles 2 and 102. Fig. 2 shows a sample DCIR pulse sequence.
@@
    $\Delta Q_{\text{DCIR}}(V)$ Features — Recognizing that each current pulse from the DCIR measurements from partial $Q(V)$ curves, the methodology from Severson et al. (2019) was extended to calculate the same set of univariate time-series statistics (mean, minimum, maximum, variance, skewness, and kurtosis) on these partial curves. Considering partial $Q(V)$ curves across the 10 SOC levels, between cycles 2 and 102, generated 60 additional features, which are referred to as $\Delta Q^{(k)}_{\text{DCIR}}(V)$ features, where $k$ represents the pulse number ($k=0$ is the pulse at 100 % SOC). This approach enriches the feature set by capturing nuances in voltage response at various SOC levels.
@@
    Linear Fit Features — To capture the overall trend in discharge capacity over time, a linear fit is applied to the discharge capacity vs. cycle number curve between two early cycles (e.g., cycles 30 and 99). The slope ($p_{1}$) and intercept ($p_{2}$) of this linear fit were used as features. These features provide insights into the rate of capacity fade and the initial state of the battery. Although more complex fits were explored, they did not significantly improve model performance and were therefore not included in the final feature set.
@@
These features were combined to train various machine learning models. For regression-based modeling, the best performance was achieved using a simple linear Elastic Net model. While more complex models were tested (as detailed in the Appendix), the Elastic Net model was selected due to its regularization properties, which help mitigate overfitting, especially when using a relatively small dataset or when the features exhibit multi-collinearity. This is particularly important given the nature of our feature sets, which include correlated variables from both traditional voltage-capacity profile features and those derived from DCIR measurements. Importantly, as a pre-processing step, Sequential Feature Selection (SFS) was applied to the high dimensional feature set to refine the input features and reduce the dimensionality of the feature space. The resulting feature space dimension was dependent on the cross-validation scheme (discussed in the following paragraphs). Principal Component Analysis (PCA) was then applied to the dimensionally-reduced feature sets. This method demonstrates the robustness of the selected features which achieve high prediction accuracy without the need for complex or resource-intensive models. Although the models trained in this work show strong predictive accuracy, it is noted that a larger and more diverse dataset is essential to further validate these findings and ensure broader applicability. While cross-validation provides a robust estimate of model performance, the addition of an independent holdout dataset would offer a more definitive evaluation of the model’s capacity to generalize across different chemistries, manufacturers, and operating conditions. Due to the limited dataset size, particularly considering the use of triplicate cells, a holdout set was not deemed feasible for this work. Instead, various cross-validation strategies were considered to assess the generalizability of the feature set.
3.2 Cross-validation strategies
To evaluate the generalization performance of the trained models, three distinct cross-validation strategies were employed, each focusing on different aspects of the dataset variability.